OpenStudy (anonymous):

The random variable X is normally distributed with mean 82 and standard deviation 7.4. Find the value of q such that P(82-q

OpenStudy (anonymous):

\[\begin{align*}P\left(82-q<X<82+q\right)&=P\left(\frac{82-q-82}{7.4}<\frac{X-82}{7.4}<\frac{82+q-82}{7.4}\right)\\ &=P\left(-\frac{q}{7.4}<Z<\frac{q}{7.4}\right)\\ &=2P\left(0<Z<\frac{q}{7.4}\right)\\ &=2\left[P\left(Z<\frac{q}{7.4}\right)-\frac{1}{2}\right]\\ 0.44&=2P\left(Z<\frac{q}{7.4}\right)-1\\ 1.44&=2P\left(Z<\frac{q}{7.4}\right)\\ 0.72&=P\left(Z<\frac{q}{7.4}\right) \end{align*}\] The probability occurs for a \(z\) value of approximately \(0.583\), so you must have \[\frac{q}{7.4}=0.583~~\iff~~q=4.314\]